Greenhab Report – April 20th

GreenHab Report
Name: Andreea Radulescu
Date: 20 Apr 2018

Environmental control:

Ambient with door open but only for few hours in the afternoon Shade cloth: on
Fan: not on.

Average temperatures: 30℃
Low 13.6℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 9 gal

Time(s) of watering for crops: Twice: In the morning around 8:45 am and afternoon 5:30pm.

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: No significant changes. Picked another batch of tomatoes as well as a couple of peppers.

Support/supplies needed: None

EVA Report – April 20th

EVA Report

EVA #18

20 Apr 2018 – Sol 12

Crew members: Richard Blakeman, Ashok Narayanamoorthi

Objectives: Exploratory reconnaissance of Cactus Road (Route 1104), continued testing of helmet fogging study and hydration pack.


Narrative: Conducted an exploratory reconnaissance of Cactus Road (Route 1104) via Galileo Road (Route 1104). Stopped at interesting geologic formations for detailed examination and to collect video and photo imagery. Conduct both mounted and dismounted EVA.

Blue ATVs #1 and #2


EVA #19

20 April 2018 – Sol 12

Crew Members: Richard Blakeman, Ashok Narayanamoorthi

Objectives: Exploratory reconnaissance of Reservoir Dam, Tank Wash, and Factory Butte View. Continued testing of helmet fogging study and hydration pack.


Narrative: Conducted an exploratory reconnaissance of the Reservoir Dam, Tank Wash, and Factory Butte View regions. The EVA was conducted using a combination of ATV and dismounted scouting. Stopped to examine multiple geologic formations and collected video and photo imagery.

Blue ATVs #3 and #4

Sol Summary – April 20th

Crew 192 Sol Summary Report 20Apr2018

Sol 12

Summary Title: Splash Down

Author: Victoria Varone

Mission Status: Nominal

Sol Activity Summary: The crew divided up our last day in sim into two fantastic EVAs and a mission to clean the hab from top to bottom for the incoming crew and close out all of our reports. We went through each of the buildings in the hab site and made sure everything is clean, organized, and in great shape for use by future crews.

Look Ahead Plan: Finishing up our last reports and inventories to prepare for incoming crew and our departure.

Anomalies in Work: None

Weather: Cloudy and rainy in the morning, warmer and sunny in the afternoon.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, Photo Report, Commander’s Report, GreenHab Report, Astronomy Report, End of Mission Report, Science Report

Support Requested: Note: Filter for loft water tank is already light brown/orange in color.

Maintenance request: Lock for Musk Observatory is no good anymore, it was severely jammed and we tried various fixes but it still sticks. Recommend replacement. Also we will need fluorescent light bulbs soon, They are flickering. Finally had to secure the water under the sink in the upstairs sink because the leak has gotten worse.

Operations Report – April 20th

Crew 192 Operations Report 20Apr2018

SOL 12

Name of person filing report: Richard Blakeman

Non-nominal systems: Upper deck faucet, hot water now has to be turned off when not in use because it leaks too much, causing water loss and constant noise from the pump.

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 60% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 25% gas left in one can.

Water (trailer) – Empty water tank secure on Trailer.
Water (static) – 550 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133825.0

Toilet tank emptied: Yes. Last emptied today, April 20.

ATV’s Used: Blue ATVs 1, 2, 3, and 4

Oil Added? No but checked and OK last time road.

ATV Fuel Used: 1, 2, and 3 at 75%, 4 at 80%

# Hours the ATVs were Used today: 6

Notes on ATVs: ATVs fine, oil checked.

Deimos rover used: No
Hours: 122.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 30.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 27.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Cleaned the hab from top to bottom to prepare for incoming crew.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, two one-piece suits used, radios working fine. Comms headpieces all cleaned and prepared for incoming crew.

Summary of Hab operations: Cleaning and organizing for Crew 193.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: Last cleaning check.

Summary of RAM operations: None

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: None

Astronomy Report – April 20th

Astronomy Report

Name: Victoria Varone

Crew: 192

Date: 20 April 2018

Sky Conditions: Clear

Wind Conditions: Calm

Observation Start Time: 1:00 PM

Observation End Time: 1:30 PM

Summary: The Sun came out after a bit of rain and a lot of clouds this morning, so I took advantage of the clear skies to get in some solar observing.

Objects Viewed: Sun

Problems Encountered: None

Science Report – April 20th

MDRS Crew 192 Final Science Report (We science the sh*t out of this!)
Richard Blakeman, Executive Officer
April 20, 2018 (Sol 12)

MDRS Crew 192 Mission Science report summary:
This report describes the outcomes of the science and research projects conducted during the MDRS Blue Crew 192 mission. It should be noted that on some of the research studies, the primary data collected requires additional time for analysis and conclusions.

Spacesuit visor fogging study
This research was conducted using a double-blind study to test off the shelf cleaning products (Joy dishwashing liquid and Johnson’s Baby shampoo) and their effectiveness against spacesuit visor fogging. Both one-piece and two-piece (separate helmet) spacesuit configurations were tested along with random controls to identify variables and collect data. The data will require analysis before any final conclusions can be made; however, preliminary data suggests that exertion levels contribute to fogging phenomenon. Additionally, baby shampoo appears to have slightly superior results in fog reduction. There were occasional reports of minor irritation but it appears to be not significant.

Hand exercises using hand relief, well-being balls
This research was conducted as single blind study to test the use of well-being ball for had exercise before EVA determining the dexterity and comfort of hands. After few measurements, discontinued the study as the exercises were creating discomfort for the crew

Crew wellness observations This is survey-based study using the Wellbeing questionnaire before, during and the end of the study to measure the happiness scale of the crew

Crew weight measurements and analysis (EVA) Daily weight measurements were taken along with the pre and post EVA analysis. Preliminary results indicated the weight loss after EVA is proportional with duration of EVA and physical exertion

Crew muscle measurements Daily crew skeletal measurements including deltoid and calf muscles were taken. Preliminary analysis show reduction in deltoid muscle in majority of the crew.

MAG (Maximum Absorbency Garment) utilization (MAG, i.e. Depend-type undergarments) were worn by crewmembers on all EVAs. This provided additional crew comfort, health, welfare, and safety protection on increasingly longer and complex EVAs. NASA and the military use the MAG protocol for missions involving extended operations involving pressure suits, EVA space suit, and undersea hard suits where waste collection issues can significantly impact crew heath and mission success.

Ultrasonic rodent repulsion experiment Three off the shelf plug-in ultrasonic rodent repulsion emitters were placed in the lower habitat, crew quarters deck, and the upper level deck. There was only one intrusion of a rodent during the mission located on the crew living deck near the refrigerator. A trap was baited with a small piece of bread coated with peanut butter and the intruding rodent was captured unharmed. On a subsequent EVA the rodent was released on Galileo Road (Route 1104). An additional rodent intruder was discovered during the night in the south-side, upper level, interface between the wall and the habitat roof structure. The intruder rodent was caught in a glue trap and did not survive. The initial conclusion is that the ultrasonic rodent repulsion emitters are ineffective. Physical traps need to be continuously deployed to capture intruder rodents and additional repulsion technologies tested.

Astronomy discussions and visual observations Conducted night time observational astronomy lectures describing various constellations and planets. The crew was able to observe several satellites and wonder at the incredible view of the heavens above. Additionally, conducted daytime solar observations using the MDRS solar telescope array. However, computer interface issues and some clouds affected viewing. Some imagery was obtained using the optical sun lens and a smart phone.

Geology observations conducted during EVAs Each EVA offered a rich and immersive experience into the local geology. Close physical inspection of structures as well as photographic and video imagery was taken for later discussion and analysis.

EVA touch screen glove testing The crew brought several types of touch pad sensitive gloves to use during EVAs. These proved to be an invaluable tool for the crews as it allowed direct interface with multiple electronic recording devices. Recommend that these be used by future crew to assist with video and photographic imagery.

Water contamination prevention and mitigation procedures
All of the habitat water storage tanks were meticulously cleaned and sanitized over the course of many days to remove any contamination and tank residue; additionally, multiple fresh water transport and loading runs to and from Hanksville was accomplished by the crew. The water transfer pump was also meticulously clean to prevent future contamination. The main water filter was also replaced by the crew.

Yuri’s night distilled spirits experiment The distillation of a celebratory spirit was both a crew morale booster and a fascinating chemistry experiment. The process took several days to complete and the resulting product was equally distributed to each crewmember in a celebratory toast to the accomplishments of Cosmonaut Yuri Gagarin for becoming the first human in space April 12, 1961.

Still and video imagery Still and video imagery has been collected by all crewmembers throughout the mission in order learn and better appreciate the challenges and requirements necessary to be an effective Martian crew.

Spacesuit hydration prototype system operational testing and evaluation An experimental prototype EVA hydration system was constructed and operationally tested on multiple EVAs both mounted and dismounted. This system has shown promising results as it can be utilized while operating a rover, ATV, as well as dismounted EVAs. Astronaut hydration, particularly during heavy exertion, is an important physiological need and critical to crew safety.

Warm Regards from Florida,

Joseph Dituri, PhD (c), CDR, US Navy Saturation Diving Officer (ret)

Commander Report – April 20th

Joseph Dituri
20 April 18

Woke to a great breakfast and a happy crew….hapoy to both clean and do final EVAs. During EVA crew left behind … ALL field day-ed lower and upper HAB inside moving all furniture.

Despite having been exceptionally clean within the HAB…another pack rat showed up. Verified not the same one as before although simular. It was caught in a glue trap on 3rd upper deck. It did not survive the ordeal and was disposed of.

Dusted all flat surfaces. Organized shelves near bathroom and downstairs and cleaned what could be cleaned.

Great geology and views. Returned from second (and final) exploration EVA and had pizza and pasta for the last supper in SIM.

IMPRESSED We are impressed with PHEnOM mission support Jamie. Thank you for getting us here! BZ Jamie and PHEnOM.

Crew intends to depart area at 0800 on Sunday.

Thank you for the opportunity to learn! Last Commander’s report.

~Dituri sends

Warm Regards,

Joseph Dituri, PhD(c)
CDR, U.S. Navy Diving Officer (ret)

Director – IBUM

Copyright © The Mars Society. All rights reserved. | Main Site